OpenStudy (anonymous):

dervative of y=4cos^5(3sqrt6x^2)

OpenStudy (anonymous):

@Zarkon

OpenStudy (zale101):

\(y=4cos^5(3sqrt6x^2)\) \(y=4(cos(3\sqrt{6x^2})^5)\) Factor out the constant and take the derivative of the rest by using the chain rule.

OpenStudy (zale101):